### v1

# Automated Statistical Model Discovery with Language Models

Abstract
Statistical model discovery involves a challenging search over a vast space of models subject to domain-specific modeling constraints. Efficiently searching over this space requires human expertise in modeling and the problem domain. Motivated by the domain knowledge and programming capabilities of large language models (LMs), we introduce a method for language model driven automated statistical model discovery. We cast our automated procedure within the framework of Box’s Loop: the LM iterates between proposing statistical models represented as probabilistic programs, acting as a modeler, and critiquing those models, acting as a domain expert. By leveraging LMs, we do not have to define a domain-specific language of models or design a handcrafted search procedure, key restrictions of previous systems. We evaluate our method in three common settings in probabilistic modeling: searching within a restricted space of models, searching over an open-ended space, and improving classic models under natural language constraints (e.g., this model should be interpretable to an ecologist). Our method matches the performance of previous systems, identifies models on par with human expert designed models, and extends classic models in interpretable ways. Our results highlight the promise of LM driven model discovery.
1 Introduction
Modeling, or generating a parsimonious but explanatory representation of a complex system, is at the heart of scientific discovery. Model discovery is challenging because it involves searching over a vast space of candidate models subject to domain-specific constraints (e.g., find the best model that remains interpretable to domain experts). Efficiently searching over the space requires extensive human expertise: modelers need broad knowledge of different modeling approaches and must work closely with domain experts to adapt these approaches to a given problem domain. As a concrete example, consider modeling blood-glucose dynamics in Type 1 diabetes (T1D) patients; accurately modeling these dynamics can enable better insulin regulation and reduce complications from the disease. To model these dynamics, modelers need to understand biomedical models that capture blood-glucose dynamics in idealized, lab settings but they also need to understand techniques for adapting these models to handle real data with noise and missingness (Miller et al., 2020). Domain experts play a crucial role in this process: modelers must work closely with clinicians to ensure that the model is consistent with human physiology. As this example illustrates, model discovery can require significant human expertise. Automating this process could accelerate and democratize scientific discovery.
Figure 1: Language model driven automated model discovery. 1) The prompt for the LM contains the dataset in textual (or visual) form, (optionally) dataset metadata (e.g., dataset description), the code for previous probabilistic programs, and natural language feedback. 2) Given this, the LM proposes new models expressed as probabilistic programs. 3) We fit these programs using automated inference procedures provided by a probabilistic programming language. 4) After we fit models, we compute model criticism statistics that capture limitations of a modeling approach. We provide these statistics to a LM which produces natural language feedback to guide future proposals. 5) We propagate the best programs, their model criticism statistics, and natural language feedback, by inserting this information into the prompt.
Automated model discovery is not a new ambition. Previous systems have been successfully deployed for discovering physical laws (Bongard & Lipson, 2007; McKinney et al., 2006; Linka et al., 2023), reverse-engineering non-linear dynamical systems (Schmidt & Lipson, 2009), nonparametric regression (Duvenaud et al., 2013) and unsupervised learning (Grosse, 2014). However, in these systems, a human expert had to carefully design a domain specific language (DSL) of models and specify a hand-crafted search procedure for composing models in that DSL. For example, in the Automated Statistician (Duvenaud et al., 2013; Lloyd et al., 2014), a system for time series modeling, human experts defined a DSL of Gaussian process kernels and a search procedure that composes kernels via addition and multiplication. Defining the DSL and the operators for composing models requires significant modeling expertise. These systems also compromised flexibility for automation: rather than choosing a model class best-suited for a problem, experts chose models that compose conveniently. This breaks the core principle of separating modeling from inference (van de Meent et al., 2021).
As we saw above, there are two key roles in the model discovery pipeline: the domain expert and the modeler. We hypothesize that LMs can supplement these roles and reduce the human expertise required in model discovery. Our hypothesis is motivated by recent work exploring LM capabilities. First, LMs have been successfully applied to domains including law (Bommarito & Katz, 2022), medicine (Lee et al., 2023), and mathematics (Wu et al., 2023). This suggests that LMs have broad domain knowledge which may enable them to supplement the domain expert. Second, LMs can reliably write code (Rozière et al., 2023; Chen et al., 2021) which means we do not require a human expert to define a DSL. Instead, we can search over a more open-ended space of models, provided they can be expressed in a generic programming language like Python. Third, LMs have strong inductive reasoning capabilities (e.g., they can generate hypotheses from limited data) (Wang et al., 2024; Qiu et al., 2024). We hypothesize these capabilities may enable them to reason about data (Zhong et al., 2023). These capabilities, in addition to their knowledge of various modeling approaches, may enable the LM to supplement the modeler.
Leveraging LMs for automated model discovery is enticing at a conceptual level. However, in order to deliver on this promising idea, we need to make several important design choices. These design choices should exploit the strengths of LMs but remain grounded in principled statistical modeling. First, we need a representation of a statistical model that can be expressed programmatically and is amenable to automated inference (e.g., model fitting). Second, we need a method for guiding LM proposals through natural language.
Algorithm 1 Automated Model Discovery with LLMs
0:  dataset $\mathcal{D}$, number of rounds $T$, $k$ number of exemplars, $m$ number of proposals per round, (optional) warm-start example $z_{0}$, function for scoring a program score, (optional) function for producing natural language feedback synthesize
  $\mathcal{Z}\leftarrow\emptyset$
  while $t<T$ do
     $\{z_{t}\}_{i=1}^{m}\sim q_{\text{LM}}(\cdot|\mathcal{Z},z_{0},h^{t-1},\mathcal{D})$
     $\{s_{t}\}_{i=1}^{m}\leftarrow\texttt{score-all}(\texttt{score},\{z_{t}\}_{i=1}^{m},\mathcal{D})$
     $\mathcal{Z}\leftarrow\texttt{select-exemplars}(k,\{z_{t}\}_{i=1}^{k},\{s_{t}\}_{i=1}^{m})$
     $h_{t}\leftarrow\texttt{synthesize}(k,\{z_{t}\}_{i=1}^{k},\{s_{t}\}_{i=1}^{k})$
  end while
To fulfill these requirements, we draw on research in probabilistic programming and inductive reasoning with LMs. In particular, we introduce the following method: LMs propose statistical models expressed as probabilistic programs, given a dataset and some metadata (e.g., dataset description). We then fit these models using generic probabilistic inference techniques, compute statistics assessing the model fit, summarize findings from these statistics in natural language, and select exemplar models to guide the next round of proposals (Figure 1). Our approach is connected to recent work on hypothesis search and inductive reasoning with LMs, but targets a fundamentally different problem (Wang et al., 2024; Qiu et al., 2024). While we also leverage the inductive reasoning capabilities of LMs, our focus is on statistical modeling of noisy real-world data, while previous work focused on learning deterministic input-output rules that can be implemented with standard Python programs.
We evaluate our method in three settings that cover common use cases in probabilistic modeling: searching within a DSL, searching over an open-ended space of probabilistic models, and improving expert models subject to modeling constraints expressed in natural language. In the first use case, we illustrate that our LM system is effective even in a DSL and matches the performance of the Automated Statistician (Duvenaud et al., 2013). We then consider the more general setting of automatically constructing probabilistic models for real world data; crucially, we do not require a user to define a DSL and this generality is enabled by our choice to use probabilistic programs. Our method identifies probabilistic programs that match the performance of human expert written programs. In the third setting, we use LMs to improve classic models and illustrate a compelling advantage of using LMs for model discovery: we can express modeling constraints in natural language (e.g., this model should be more physical) that guide LMs towards models that balance interpretability and flexibility.
2 Box’s Loop with Language Models
2.1 Background
Before introducing our method, we provide a brief overview of the probabilistic modeling paradigm. We then motivate a key design choice: representing probabilistic models with probabilistic programs. In probabilistic modeling, our goal is to describe a dataset in terms of unobserved, latent structure. Formally, a probabilistic model is a joint distribution $p(x,z|\eta);$ where $x=x_{1:N}$ denotes $N$ observed data points, $z=z_{1:M}$ denotes $M$ latent variables, and $\eta$ corresponds to non-random quantities in the model (Blei, 2014). After specifying a probabilistic model, we fit the model to observed data. Fitting a model involves inference, or computing the posterior distribution $p(z|x)$. In the model criticism step, we evaluate the model by interrogating the posterior. In this work, we use a common model criticism technique known as a posterior predictive check: to identify discrepancies, we sample from the posterior predictive distribution and compare these samples against observed data (Gelman et al., 2013). Model building and model criticism typically take place over multiple iterations in a process known as Box’s Loop (Box & Hunter, 1962).
There are many different representations of a probabilistic model. Since our goal is automated model discovery, we use probabilistic programs. Probabilistic programming languages provide ways to flexibly represent probabilistic models as programs and support generic inference methods for any arbitrary program (Wood et al., 2014; Goodman et al., 2008; van de Meent et al., 2021). In the context of automated model discovery, these are highly desirable properties, since LMs can write code reliably and other representations of probabilistic models can require custom inference methods. Our approach of using LMs to generate probabilistic programs is related to the approach taken by Wong et al. (2023) but is motivated by a different problem. Our emphasis is on using LMs as a tool for developing statistical models of real-world datasets, while their focus was on integrating symbolic methods with LMs to enhance their reasoning capabilities.
2.2 Problem formulation
We formally describe our framework. Our framework is motivated by recent work in inductive reasoning with LMs (Qiu et al., 2024; Wang et al., 2024), integrating tools with LMs (Gao et al., 2023), and driving LMs via linguistic feedback (Shinn et al., 2023).
At a high level, we consider a method for learning probabilistic models from data that involves two steps: a proposal step and a criticism step [^1] To avoid overloading the word “model”, we will refer to language models as LMs.. Crucially, by learning a model, we mean searching over the space of model structures and not just learning the parameters of some fixed model class. In each step, we leverage LMs. In the proposal step, a proposal LM proposes probabilistic programs for a dataset. We then fit these probabilistic programs and evaluate them. In the criticism step, we provide the critic LLM with programs and criticism statistics and ask the critic LLM to provide insights that guide the next round of proposals.
We start with a dataset $\mathcal{D}=\{\mathbf{x_{i}},y_{i}\}_{i=1}^{n}$. Here $\mathbf{x_{i}}\in\mathbb{R}^{d}$ are fixed $d$-dimensional input values (e.g., features) and $y_{i}\in\mathbb{R}$ are the observations. Let $\Sigma$ be the vocabulary of the LM. For each dataset, we have an associated metadata set $\mathcal{C}\in\Sigma^{d+1}$, which consists of natural language descriptions of $\mathcal{D}$ (e.g., animal ages vs length) and natural language descriptions of each feature in $\mathcal{D}$ (e.g., length of animal). Context informs how human modelers leverage prior knowledge; for example, if a modeler knows their dataset consists of monthly carbon dioxide measurements over a fifty-year time span, they will choose a model that can capture periodicity and linearity. Our goal is to find a probabilistic program $z\sim\Sigma^{*}$ that maximizes some notion of quality. We quantify this with the log marginal likelihood or expected log predictive density (ELPD) estimated via cross validation (LOO) (Vehtari et al., 2017).
2.3 Approach
We summarize our approach in Figure 1 and Algorithm 7.
Model Building Step
To generate candidate models for round $t$, we sample $m$ probabilistic programs $z^{t}_{i}$ from the proposal LM $q_{\text{LM}}(\cdot)$. In our experiments, we use GPT-4 V (Achiam et al., 2023), which has multimodal capabilities. We leverage in-context learning, or LMs ability to learn from examples in a prompt, to guide the LM’s proposals based on high-scoring programs in the past (Brown et al., 2020). The proposal LM $q_{\text{LM}}$ “conditions” on $h^{t-1}\in\Sigma^{*}$, a natural language instruction synthesizing previous modeling approaches and suggesting new approaches, $k$ exemplars $\{z_{1},\ldots z_{k}\}$, and a visual or textual representation of $\mathcal{D}$. In later analyses, the proposal LM also conditions on a warm-start expert program $z_{0}$.
$z_{i}^{t}\sim q_{\text{LM}}(\cdot|z_{0},z_{1},\ldots,z_{k},h^{t-1},\mathcal{D})$
Chain-of-thought reasoning, or generating intermediate reasoning steps, improves the performance of LMs (Wei et al., 2022; Kojima et al., 2022). Motivated by this, we instruct the proposal LM to reflect on the properties of the dataset in the plot of the data, sketch a high-level modeling approach, state the hypotheses that it will address before writing a program, and add comments to code that address specific hypotheses.
To create the round $t$ exemplars for the Proposal LM, we choose the best $k$ programs among the $m$ proposed programs in round $t-1$. Alternatively, we could choose the $k$ best programs across all rounds. However, we sometimes noticed that this alternative approach suffers from a pathology: all future proposals might collapse to a single “ancestor” program from an early round. This can happen if a better model class is proposed in round $t$ (e.g., hierarchical model) but the specific parameterization under-performs a model proposed earlier (e.g., non-hierarchical).
Scoring Step
In the scoring step, we fit a probabilistic program to data. Fitting this model requires approximate inference so we leverage pymc (Abril-Pla et al., 2023), a probabilistic programming library in Python. pymc automatically assigns a Markov Chain Monte Carlo (MCMC) sampler to perform inference. Crucially, by using a probabilistic program, we decouple modeling from inference: the proposal LM focuses on constructing the best model, while pymc takes care of inference. Our approach of offloading computations to an external tool is connected to recent work augmenting LM capabilities through interaction with tools (e.g., Python interpreter) (Gao et al., 2023; Schick et al., 2023).
Model Criticism Step
In the criticism step, we ask the LM to identify weaknesses in the models proposed so far. To help the LM identify weaknesses, we compute model criticism statistics for a list of previously proposed probabilistic programs. We then provide these model criticism statistics to a critic LM to identify what modeling approaches worked. Naively, we could simply provide all proposed programs to the critic LM. However, this list grows each round and potentially has redundancy. Furthermore, LMs struggle to reason over long contexts. On the other hand, only providing the top programs in a given round might overly bias the LM to the most recent round of proposals. Since probabilistic programs can sometimes be summarized succinctly in natural language (e.g., von Bertalanffy growth function with informative priors), we distill information about previous approaches by asking a critic LM to provide linguistic feedback $h^{t-1}$ on programs (Shinn et al., 2023). In particular, to produce $h^{t-1}$, we sample from a critic LLM $p_{LM}$ and again leverage in-context learning. Let $\tilde{z}_{1},\ldots\tilde{z}_{d}$ be the top $d$ programs across all rounds of revision. We generate $h^{t-1}$ as follows
$h^{t-1}\sim p_{\text{LM}}(\cdot|\tilde{z_{1}},\ldots,\tilde{z_{k}},h,\mathcal{D})$
Here, $h$ is an instruction to the LM for what kind of feedback to provide. In our experiments, we ask the LM to 1) provide natural language hypotheses for discrepancies between the model and the data and 2) provide a synthesis of what strategies/approaches worked. Crucially, to capture long-term dependencies and reduce bias towards the recent round, the programs $\{\tilde{z_{i}}\}_{i=1}^{d}$ provided to the critic LM are the best $d$ programs across all rounds; in general, this need not overlap with the exemplars $\{z_{j}\}_{j=1}^{m}$ provided to the proposal LM.
3 Experiments
3.1 Searching over a DSL: automated Gaussian process kernel discovery
Dataset LM+ LM Periodic AS SM N-BEATS
Air 0.08 0.07 0.15 0.19 0.06 0.22
Beer 0.07 0.22 0.15 0.06 0.05 0.02
Heart 0.21 0.21 0.20 0.21 0.21 0.07
Milk 0.12 0.10 0.10 0.11 0.09 0.04
Wine 0.14 0.16 0.21 0.13 0.18 0.17
Wool 0.20 0.19 0.19 0.23 0.13 0.18
Figure 2: Test set performance of LM system on time series datasets. The LM system identifies compositional kernels that match the performance of strong baselines, such as the Automated Statistician. (left) Comparison of Test Mean Absolute Error (MAE) against Automated Statistician (Greedy Search), Spectral Mixture kernel, Periodic kernel, and N-BEATS. LM+ refers to search over augmented base kernel space. We bold the best and underline second among GP methods. (right) Extrapolations from GP with LM discovered kernel.
We first evaluate LM’s ability to search over a constrained space of models; in some settings, a domain expert may require the modeler to search over a DSL for reasons such as interpretability. In particular, we consider time-series modeling with Gaussian processes (GPs) as in the Automated Statistician (Duvenaud et al., 2013). GPs are probabilistic models that specify a distribution over functions. GPs are defined by a mean function and a positive-definite kernel function $k(\mathbf{x},\ \mathbf{x^{\prime}})$ that gives the covariance between $f(\mathbf{x})$ and $f(\mathbf{x^{\prime}})$ as a function of $\mathbf{x}$ and $\mathbf{x^{\prime}}$. The properties of a GP depend significantly on the kernel and therefore choosing a kernel that reflects domain knowledge (e.g., linearity, periodicity) is a key design choice. One common way to produce a more flexible kernel is to compose kernels via addition and multiplication, leveraging the closure properties of kernels. Here, we evaluate LMs ability to search over a space of kernels to identify an appropriate composition of kernels.
Setup
Mirroring Duvenaud et al. (2013), we define a space of base kernels. In the prompt, we ask LM to perform one of the three operations (addition, multiplication, and replacement) on one of the in-context exemplar programs. For the first round, we ask the LM to produce an initial guess based on the structure of the dataset, which we provide as a plot. Given a kernel expression, we learn the kernel hyperparameters via gradient-based optimization of the marginal likelihood.
Results
We evaluate our method on six common univariate time series datasets. We compare against the Automated Statistician, a greedy algorithm proposed by Duvenaud et al. (2013) (run until a depth of 10), and two GP baselines: a periodic kernel, a spectral mixture kernel (a strong non-compositional baseline). We also compare against N-BEATS, a strong neural baseline (Wilson & Adams, 2013; Oreshkin et al., 2019). In Figure 2, we compare the mean absolute error (MAE) on held-out test data for all datasets. To account for stochasticity in the Automated Statistician implementation we used (Saad et al., 2023) and stochasticity in different repetitions of our pipeline, we average the test MAE across three model runs. Our method, denoted LM in the table, matches the performance of the Automated Statistician, showing that the LM can efficiently search over a constrained space of models. We also experiment with augmenting the base kernel space with additional kernels (denoted LM+ in the table). In some cases, the LM can leverage this additional flexibility; for example, by using the additional kernels, LM+ can better capture the Australian beer sales dataset than LM. On the right panel, we show the extrapolations and identified kernel for the monthly air passenger dataset; LM+ identifies a kernel with a periodic times linear component to capture the increasing amplitude in the data.
3.2 Open-ended probabilistic model discovery for real-world datasets
Dataset Expert LM
Eight schools -30.70 -30.42
Eight schools sim -18.09 -18.31
Eight schools sim no metadata -18.09 -16.36
Dugongs 22.52 23.40
Dugongs sim 50.04 57.40
Dugongs sim no metadata 50.04 55.24
Surgical -40.29 -38.03
Surgical sim -39.80 -38.38
Surgical sim no metadata -39.80 -63.72
Peregrine -142.19 -173.11
Peregrine sim -130.48 -179.06
Peregrine sim no meta -130.48 -136.39
Table 1: Comparison of LM programs against expert programs We perform this comparison across four different datasets and two different ablations that replace observations with synthetic observations and remove all metadata. We report the expected predictive log density estimated via leave-one-out cross validation. We bold statistically significant differences and underline non-significant differences. LM programs match the performance of human expert programs on 9/12 datasets.
By integrating LMs into the model discovery process, we can search over a much broader class of models. Here, we explore LMs ability to automatically construct probabilistic programs for datasets. We prompt a LM to produce probabilistic programs in pymc, a probabilistic programming language in Python, enabling us to search over a vast space of statistical models (Abril-Pla et al., 2023).
Dataset
We consider four real world datasets from the Stan PosteriorDB dataset (Magnusson et al., 2023): a dataset consisting of average improvement in SAT scores after a SAT improvement program across eight different high schools, a dataset consisting of ages of twenty-seven dugongs and their lengths, a surgical dataset consisting of mortality rates in twelve hospitals performing cardiac surgery on babies, and a dataset of peregrine population counts in the French Jura from 1964 to 2003. Each dataset has an associated human expert written probabilistic program in Stan, which we translate into pymc. These expert programs are generally open-source contributions from the Stan developer community. These datasets cover common modeling motifs such as hierarchical modeling and regression.
Ablations
To study how domain knowledge affects the LM’s modeling capabilities and to mitigate concerns about dataset leakage, we consider two ablations. In the first ablation, we construct a simulated analog for each dataset: for each dataset, we generate synthetic observations by sampling from the prior of a generative model or by sampling from a model with fixed parameters. We denote these datasets as simulated datasets. For example, to generate a simulated dataset for the eight schools dataset, we sample from the prior of the human expert program. To generate a simulated analog of the dugongs dataset, we fix the values of the parameters $\alpha,\beta,\gamma$ used in the expert model $y_{i}=\alpha-\beta\gamma^{x_{i}}$. Since it is highly unlikely that these simulated datasets appeared in GPT-4’s training data, this ablation helps mitigate concerns about dataset leakage.
In the second ablation, we remove dataset metadata from the information provided to the LM. In particular, we remove the dataset description, replace the column names with domain-agnostic column names (e.g., $x_{0},y$) and replace the axis labels with uninformative labels. We denote these datasets as no metadata datasets. By removing this metadata, we can characterize how LMs use domain knowledge and mitigate concerns about dataset leakage.
Quantitative Results
We compare the performance of our method across different datasets and ablations against expert programs in Table 1. We emphasize that the expert programs are strong baselines especially for the simulated datasets since the expert programs generated the data. We highlight numbers corresponding to significant differences, where significance is defined as an ELPD difference of greater than four times the standard error estimate.
LMs reliably identify programs on par with expert programs; we validate the fidelity of the posterior distributions using standard Markov Chain Monte Carlo diagnostics. Crucially, removing metadata or switching to simulated datasets does not generally reduce performance relative to the expert program. The main exception is the surgical dataset. By replacing the labels with $x_{0}$ and $y$, the LM formulates this problem as a regression problem instead of using a hierarchical model like the expert model. For the Peregrine dataset, removing metadata improves performance. We discuss this further in the next section.
Qualitative Analysis: Language Models are Domain Experts
Earlier, we asked whether LMs can reliably play the role of a domain expert. In our ablations, we study how metadata influences the model search process by examining how removing metadata changes the programs proposed by LMs. In Figure 3, we plot the model fit and list the corresponding LM programs in the figure caption. For the peregrine population dataset, when given all the metadata, the LM models this dataset using a logistic growth model, which is motivated by the problem domain. However, logistic growth models cannot capture the initial decline in population. Interestingly, when we remove all metadata, the LM performs better, because it uses a quartic polynomial to model the data. This closes the gap in performance with the expert program. We observe a similar trend with the dugongs growth curve dataset. When given metadata, the LM uses a von Bertalanffy growth function (von Bertalanffy, 1949), which is commonly used to model animal growth. When we remove all metadata, the LM uses a polynomial function with a logarithmically transformed value for the inputs.
Figure 3: Domain knowledge shapes LM modeling approaches. In the top row, we keep all metadata (e.g., dataset description). In the bottom row, we remove metadata that reveals information about the domain. This leads to qualitatively different approaches for three datasets; for eight schools, the LM discovers a hierarchical model even without metadata. We list the corresponding programs for these different ablations. (top row) $\frac{K}{(1+((K-P_{0})/P_{0})\exp(-rt))}$, $\text{BetaBin}(n,\alpha,\beta)$, $L_{\text{inf}}(1-\exp(-k(\text{age}-t_{0})))$; (bottom row) $a+bx_{0}+cx_{0}^{2}+dx_{0}^{3}+ex_{0}^{4}$, $\alpha+\beta x_{0}+\gamma x_{0}^{2}$, $\alpha+\beta_{1}\log x_{0}+\beta_{2}\log x_{0}^{2}+\beta_{3}\log x_{0}^{3}$.
Interestingly, the LM sets the prior parameters in these models in a data-informed way. For example, the LM sets the asymptotic length in the von Bertanlanffy function based on the observed lengths in the dataset. See code snippet in Figure 7 of the Appendix. For the eight schools dataset, the LM identifies a hierarchical model even without metadata, illustrating that model criticism can compensate for lack of domain knowledge. We illustrate the improvement round to round in Figure 4.
Figure 4: Round-to-round improvement. Model criticism leads to improvements on average, although improvement is not necessarily monotonic for a given dataset and run. (left) LOO score vs round for Stan dataset. We normalize the scores across a run of Box’s loop and average across Stan datasets for the no metadata, simulated condition. Larger is better and error bars correspond to standard error. (right) Squared error vs round for Lotka-Volterra experiment. We report the squared error averaged across three different runs, for the warm-start, no constraint condition; smaller is better.
3.3 Improving classic models under modeling constraints
Figure 5: Correcting misspecified Lotka-Volterra dynamics. LM can introduce corrections to standard Lotka-Volterra dynamics (no warm-start) and a hybrid neural approach (warm-start) that outperform several baselines. (left) LM proposed model predictions on training data and extrapolations (grey region); (right) LM models (warm-start and no-warm start) achieve lower test MAE than a neural ODE model, a hybrid ODE model with a multiplicative correction to the prey-predator dynamics, and the standard Lotka-Volterra (LV) model.
1def no_warm_start(t, y, _coeffs): 2  # Hypothesis 1: Logistic growth for prey with carrying capacity kappa 3  # Hypothesis 2: Saturation effect in predation rate with parameter psi 4  db_dt = alpha * b * (1 - b / kappa) - beta * b * c / (1 + psi * b) 5  dc_dt = -gamma * c + delta * b * c - epsilon * c**2 1def warm_start_constrained(t, y, _mlp): 2  inputs = jnp.array([b])  # Input is prey density (bunny population) 3  # Neural network learns a modifier for the handling time based on prey density 4  handling_time_modifier = _mlp(inputs) 5  # Modulate the predation rate beta in a non-linear manner according to handling time 6  db_dt = alpha * b - beta * b * c / (1 + handling_time_modifier[0] * b) 7  dc_dt = -gamma * c + delta * b * c / (1 + handling_time_modifier[0] * b)
Figure 6: LM can propose corrections to ODEs (top) In the no warm-start variation, the LM introduces corrections informed by domain knowledge of predator-prey models (carrying capacity, predation saturation) (bottom) When prompted to introduce neural networks in an interpretable way, one strategy the LM proposes is to make the handling time parameter depend non-linearly on the prey density, extending a traditional approach to modeling predation saturation.
In the previous experiment, we explored LM’s ability to identify models tabula rasa (e.g., without any initial seed model). However, in many scientific settings, we begin with a well-known model and are tasked with improving it. Here, we explore LM’s ability to improve upon a Lotka-Volterra model of predator-prey dynamics. We also explore LM’s ability to respect “soft” modeling constraints expressed in natural language (e.g., this model should be interpretable to ecologists), an important component of model discovery.
Dataset
To create our dataset, we simulate data from the following “perturbed” Lotka-Volterra dynamics
$\frac{db}{dt}$ $=\alpha b-\beta bc$ (1)
$\frac{dc}{dt}$ $=-\gamma c+\delta bc^{0.95}$ (2)
Setup
The LM is tasked with implementing a ODE model in Python using Jax (Bradbury et al., 2018). Estimating the parameters of Lotka-Volterra models via Bayesian inference is challenging. We instead learn the parameters via gradient descent which can be straightforwardly implemented using modern automatic differentiation libraries (Chen et al., 2018; Kidger, 2021).
We consider three variations: warm-start with constraints, warm-start with no constraints and no warm-start that differ in their initial seed program and the initial instructions which express modeling constraints. In all variations, we provide the LM with a scatter plot of training datapoints. In the no warm-start (WS) variation, we provide the LM with an implementation of standard Lotka-Volterra dynamics using diffrax (Kidger, 2021), a Jax based ODE library that supports learning ODE parameters via backpropagation and the predictions obtained from fitting standard Lotka-Volterra to the training data. In both warm-start (WS) variations, we provide the LM with an implementation of a hybrid Neural ODE that introduces an additive correction term to the predator dynamics, parameterized by a multilayer perceptron (MLP), and the predictions obtained from fitting this model to the training data. In the constrained warm-start variation, we ask the LM to produce a hybrid neural ODE model that is interpretable to an ecology expert. In the unconstrained warm-start variation, we provide the same seed program but we do not impose this additional constraint. Importantly, for models with both neural and physical components, we employ a two-stage learning procedure so that the neural component does not dominate the dynamics; see Appendix C for details.
Results
In Figure 5 (left), we plot the predictions obtained from integrating the learned dynamics for an ODE proposed by the LM, the training data points generated from the true dynamics, and the predictions from standard Lotka-Volterra. For all ODEs, we fit free parameters to the training data via gradient descent. The grey region indicates the extrapolation region.
The LM can significantly improve upon the standard Lotka-Volterra model by introducing corrections to the dynamics (Figure 5). The standard Lotka-Volterra model cannot capture the decreasing amplitude in the data; furthermore, there is a slight phase shift relative to the training datapoints. In contrast, the LM identifies an ODE that captures these properties and extrapolates accurately. In Figure 5 (right), we compare these programs against a neural ODE baseline, and a hybrid neural ODE baseline that incorporates a multiplicative correction (parameterized by an MLP) to the predator-prey interaction term in the predator equation. See Section C of the Appendix for details on these baselines. For LM variations, we report the average test MAE across three runs. All LM variations outperform the baselines.
In Figure 6, we present code snippets corresponding to representative programs proposed in the constrained and unconstrained variations. These snippets show how natural language constraints can guide the LM towards more flexible models that retain interpretability. For the variation with no natural language constraints, the LM takes a purely empirical approach. In particular, the LM adjusts the scaling terms on the additive correction term parameterized by an MLP (Figure 8). For the constrained variation, the LM proposes a hybrid approach integrating the neural approach in the prompt with classic models in the literature; importantly, even though the LM is asked to balance interpretability and flexibility, the LM still identifies programs that outperform the Neural ODE and standard Lotka-Volterra baselines. One approach the LM proposes is an extension of the Rosenzweig and MacArthur model with a Holling’s type II functional response (Rosenzweig & MacArthur, 1963) to allow a static parameter to depend dynamically on the prey density: the LM models the handling time, or the time a predator spends “processing” a prey, as a nonlinear function of the prey density via an MLP. These results show that natural language constraints drive LMs towards models that balance flexibility and interpretability.
4 Conclusion
We introduced a method for leveraging LMs for automated model discovery. We showed our method can identify models that perform favorably against strong baselines. We also studied how domain knowledge and natural language constraints influence the programs proposed by LMs. Altogether, our results highlight the compelling advantages of LM driven statistical model discovery. Although we focused on the static data regime in this work, a promising future direction is to explore settings where the LM designs experiments to collect more data.
5 Acknowledgements
This work was supported in part by AFOSR Grant FA9550-21-1-0397, ONR Grant N00014-22-1-2110, and an NSF Expeditions Grant, Award Number (FAIN) 1918771. EBF is a Chan Zuckerberg Biohub – San Francisco Investigator.
Appendix A Gaussian Process experiments
In the prompt, we ask the LM to use the following operations.
  1. 1.
    Replace a subexpression $\mathcal{S}$ with $\mathcal{S}+\mathcal{B}$, where $\mathcal{B}$ is any base kernel.
  2. 2.
    Replace a subexpression $\mathcal{S}$ with $\mathcal{S}x\mathcal{B}$, where $\mathcal{B}$ is any base kernel.
  3. 3.
    Any base kernel $\mathcal{B}$ may be replaced with any other base kernel family.
LM hyperparameters
We provide the LM with the following kernels: Exponentiated Quadratic, Periodic, Linear, and Polynomial. We run our pipeline for two rounds with three proposals each round. We use a temperature of 0.2 for the Proposal LM and temperature of 0.0 for the Critic LM. We use three in-context exemplars. Our Critic LM conditions on the best twelve programs so far.
In the augmented variation, we also provide the LM with the following additional kernels: Matern32, Matern52, Cosine, and the Rational Quadratic kernel. In the augumented variation, we run our pipeline for three rounds with eight proposals each round. We use a temperature of 0.7 for the Proposal LM and temperature of 0.0 for the Critic LM. We use three in-context exemplars. Our Critic LM conditions on the best twelve programs so far.
The marginal likelihood can be multimodal in the parameters of the periodic kernel. Therefore, following Duvenaud et al. (2013), if the proposed kernel has periodic components, we initialize the period at five different initial values, optimize the marginal likelihood starting from those different initializations, and choose the kernel hyperparameters with the highest marginal likelihood across those initializations.
Spectral Mixture kernel
We use a GP with a spectral mixture kernel (Wilson & Adams, 2013) with 5 mixture components. For each dataset, we randomly initialize the parameters of the mixture and choose the kernel hyperparameters with the highest log marginal likelihood across five random initializations.
Appendix B Stan Experiments
Eight Schools Dataset
This dataset consists of eight observations: the estimated treatment effect of a SAT coaching program and the standard error of the treatment effect.
Peregrine dataset
This dataset consists of peregrine population counts in the French Jura from 1964 to 2003 (40 observations in total).
Dugongs Dataset
The ages and lengths of 27 captured dugongs (sea cows).
Surgical Dataset
The mortality rates in 12 hospitals performing cardiac surgery on babies.
LM hyperparameters
We run our pipeline for three rounds with eight proposals each round. We use a temperature of 0.7 for the Proposal LM and temperature of 0.0 for the Critic LM. We use three in-context exemplars. Our Critic LM conditions on the best twelve programs so far.
Markov Chain Monte Carlo diagnostics
We evaluated the fidelity of the learned posteriors using the Gelman-Rubin $\hat{R}$ diagnostic (Gelman & Rubin, 1992) and by examining the Bulk Effective Sample Size (ESS). In particular, the programs reported in the table all had $\hat{R}\leq 1.01$ and mean bulk ESS $>=400$ per chain.
Language model produced pymc code snippet for modeling Peregrine population count
1def dugongs_gen_model():
2  # Priors for von Bertalanffy growth function parameters
3  L_inf = pm.Uniform(’L_inf’, lower=0, upper=3) # Asymptotic length, bounded by observed data
4  k = pm.Uniform(’k’, lower=0, upper=1) # Growth coefficient, bounded by reasonable values
5  t0 = pm.Uniform(’t0’, lower=-5, upper=5) # Theoretical age at zero length
6  # Expected length at each age using von Bertalanffy growth function
7  expected_length = L_inf * (1 - pm.math.exp(-k * (ages - t0)))
8  # Likelihood for observed lengths
9  sigma = pm.Uniform(’sigma’, lower=0, upper=1) # Standard deviation of observed lengths around the mean
10  y_obs = pm.Normal(’y_obs’, mu=expected_length, sigma=sigma, observed=lengths)
Figure 7: LM proposes programs informed by domain knowledge LM chooses a model informed by the domain (animal length vs age) and sets the priors based on the dataset (e.g., the largest length in the dataset is smaller than 3).
Appendix C Lotka-Volterra
Dataset
To create our dataset, we simulate data from the following “perturbed” Lotka-Volterra dynamics
$\frac{db}{dt}$ $=\alpha b-\beta bc$ (3)
$\frac{dc}{dt}$ $=-\gamma c+\delta bc^{0.95}$ (4)
We set $\alpha=0.9,\beta=1.1,\delta=-1.2,\gamma=2.1$. The parameter $\alpha$ characterizes the prey’s maximum growth rate and $\beta$ controls how the predator population modulate the growth rate. The parameter $\gamma$ characterizes the prey’s maximum death rate and $\delta$ controls how the predator’s growth rate depends on the prey population density. In contrast to the standard Lotka-Volterra dynamics, we raise $c$ to a fractional power.
1def no_warm_start(t, y, _coeffs): 2  # Hypothesis 1: Logistic growth for prey with carrying capacity kappa 3  # Hypothesis 2: Saturation effect in predation rate with parameter psi 4  db_dt = alpha * b * (1 - b / kappa) - beta * b * c / (1 + psi * b) 5  dc_dt = -gamma * c + delta * b * c - epsilon * c**2 1def warm_start_constrained(t, y, _mlp): 2  inputs = jnp.array([b])  # Input is prey density (bunny population) 3  # Neural network learns a modifier for the handling time based on prey density 4  handling_time_modifier = _mlp(inputs) 5  # Modulate the predation rate beta in a non-linear manner according to handling time 6  db_dt = alpha * b - beta * b * c / (1 + handling_time_modifier[0] * b) 7  dc_dt = -gamma * c + delta * b * c / (1 + handling_time_modifier[0] * b) 1def warm_start_unconstrained(t, y, _mlp): 2  inputs = jfu.ravel_pytree((b, c))[0] 3  mlp_output = _mlp(inputs) 4  # Fine-tuned scaling of the MLP output to match the amplitude of data more closely 5  db_dt = alpha * b - beta * b * c + 0.02 * mlp_output[0] # Reduced scaling for bunnies 6  dc_dt = -gamma * c + delta * b * c + 0.06 * mlp_output[1] # Increased scaling for cats
Figure 8: LM can propose corrections to ODEs (top) In the no warm-start variation, the LM introduces corrections informed by knowledge of predator-prey models (carrying capacity, predation saturation) (middle) When prompted to introduce neural networks in an interpretable way, LM proposes to make the handling time parameter depend non-linearly on the prey density, extending the Holling’s type II response, a traditional approach to modeling predation saturation. (bottom) When prompted to introduce neural networks without constraints, LM introduces additive MLP parameterized corrections and adjusts the scaling factors.
We now describe the various baselines we compare against in Section 3.3.
Standard Lotka-Volterra
We fit the free parameters of the standard Lotkva-Volterra differential equations.
$\frac{db}{dt}$ $=\alpha b-\beta bc$ (5)
$\frac{dc}{dt}$ $=-\gamma c+\delta bc$ (6)
Neural ODE baseline
We parameterize the predator and prey equations with an MLP. We run a hyperparameter search over four widths (4, 8, 16, 32) and 3 depths (1,2,4). We use a learning rate of 3e-3 and train using full-batch gradient descent with Adam for 1500 iterations.
Hybrid Neural ODE baseline
We implement a Hybrid Neural ODE baseline that introduces a correction to the predator-prey interaction term. Note that, we follow a two-stage “boosting” type procedure to fit the parameters of the MLP. First, we fit the free parameters $\alpha,\beta,\gamma,\delta$ to the data. We then freeze those parameters and fit the MLP parameters. Without this two-staged approach, the MLP term can dominate the dynamics. The MLP term has one layer and four hidden units and we train the MLP with full batch gradient descent with Adam using a learning rate of 3e-3.
$\frac{db}{dt}$ $=\alpha\cdot b-\beta\cdot b\cdot c$ (7)
$\frac{dc}{dt}$ $=-\gamma\cdot c+\delta\cdot b\cdot(c+0.1\cdot\text{\text{mlp}}(b,c))$ (8)
In the warm-start variations, we provide the LM with an initial hybrid Neural ODE baseline that introduces an additive correction to prey equation. The MLP term has one layer and four hidden units.
$\frac{db}{dt}$ $=\alpha\cdot b-\beta\cdot b\cdot c+0.1\cdot\text{\text{mlp}}(b,c)$ (9)
$\frac{dc}{dt}$ $=-\gamma\cdot c+\delta\cdot b\cdot c$ (10)
LM hyperparameters
We run our pipeline for four rounds with twelve proposals each round. We use a temperature of 0.7 for the Proposal LM and temperature of 0.0 for the Critic LM. We use three in-context exemplars. Our Critic LM conditions on the best twelve programs so far.